OpenStudy (anonymous):

Teesha is in the French club. There are 31 students in the club. The French teacher will pick two students at random to guide visiting students from France. What is the probability that Teesha will be picked as a guide? Hint: Teesha must be picked first or second A. 1/30 B. 30/31 C. 1/930 D. 2/31

hartnn (hartnn):

just solved the same type of question.... http://openstudy.com/users/smokeysthename#/updates/51518fd7e4b0ae0b658be4da

OpenStudy (anonymous):

Thank you hartnn, I'll take a look at that.

hartnn (hartnn):

welcome ^_^ if you got any doubts, ask here, we'll clarify it here itself.

OpenStudy (anonymous):

Thank you, will do. :b

OpenStudy (anonymous):

Mhmm Mhmm so I went and looked............ I'm extremely lost o.o lol >.<

hartnn (hartnn):

the probability that Teesha will be picked as a guide means Teesha will be picked on 1st selection AND Teesha will be picked on 2nd selection got this ?

OpenStudy (anonymous):

Yesh

hartnn (hartnn):

i would like to revise my previous statement, the probability that Teesha will be picked as a guide means Teesha will be picked on 1st selection and NOT in 2nd(because she's already selected), AND Teesha will be picked on 2nd selection and NOT in 1st.

OpenStudy (anonymous):

Oooooh, so what happens then?

hartnn (hartnn):

so, we'll consider 1st case, Teesha will be picked on 1st selection and NOT in 2nd(because she's already selected), 1st selection, teacher has 31 choices, to select teesha, she has 1 option, so, 1/31 2nd selection, teacher now have 30 other choices, and to NOT select teesha, she can pick any of 30 other, so 30/30 so, overall for 1st case, it'll be 1/31 *30/30 got this ?

OpenStudy (anonymous):

I got it...... 30/930, does it need to be simplified?

hartnn (hartnn):

1/31 + 30/31*1/30 = 2/31 how you got 30/930 ??

OpenStudy (anonymous):

Lol............ ooooooh goodness, I multiplied 1/31 *30/30 o.o

hartnn (hartnn):

ohh..for 1st case....30/30 =1 :P i though you've got the final answer.

hartnn (hartnn):

ok, did you understand 1st case ? can i move to 2nd case ?

OpenStudy (anonymous):

Yes please, moving on :b

hartnn (hartnn):

2nd case, Teesha will be picked on 2nd selection and NOT in 1st., 1st selection, teacher has 31 choices, to NOT select teesha, she has 30 other options, so, 30/31 2nd selection, teacher now have 30 other choices, and to select teesha, she have to pick her, so 1/30 so, overall for 1st case, it'll be 30/31*1/30 = 1/31 got this ?

hartnn (hartnn):

since the 2 cases are independent, you'll add them, 1/31 + 1/31 = 2/31

OpenStudy (anonymous):

Ok then

OpenStudy (anonymous):

So 2/31 * 1/31 would = 2/31 as the answer? Lol, or am I too early?

hartnn (hartnn):

1/31 + 1/31 = 2/31 thats why 2/31 is the answer as we considered both the cases.

OpenStudy (anonymous):

Ohhhhhhh lol ok, I understand that now :b

OpenStudy (anonymous):

Thank you [: You've been fanned lol.

hartnn (hartnn):

thanks for fanning! :D and welcome ^_^